# A Gaussian smooth transition vector autoregressive model: An application to the macroeconomic effects of severe weather shocks

### v2

@@ 4.2 GSTVAR model
We are predominantly interested in the economic impact of severe weather, and, therefore, a severe weather shock (ACI shock) must be identified. This amounts to imposing a unique structure of the impact matrix $B_{t}$ in (2.3) governing the contemporaneous relationships of the shocks, so that one of the shocks can be labelled the ACI shock. Following Kim et al. (2022), our key identification restriction is that the ACI shock is the only shock that can instantaneously affect the ACI index. This identification restriction seems reasonable, as it states that the other (macroeconomic) shocks do not affect severe weather or sea level within a month, but allows them to affect the ACI in the long run. We establish the identification by placing the ACI index first in the vector of variables and imposing a recursive lower-triangular structure on the impact matrix $B_{t}$, obtained by the Cholesky decomposition of the estimated conditional error covariance matrix $\Omega_{y,t}$.
4.3 Impulse response analysis
To study the effects of the ACI shock, we compute the generalized impulse responses of the variables to it. As pointed out in Section 2.2, due to the endogenously determined transition weights and the fact that we allow the regime to switch as a result of a shock, there are multiple types of possible asymmetries. Therefore, we compute generalized impulse responses that can vary depending on the initial values of the variables as well as on the sign and size of the shock. We are particularly interested in finding out about the potential state-dependence of the effects of the ACI shock, and to capture them, we draw initial values from the stationary distribution of each regime separately. Shocks of different sign and size seem to result in similar impulse responses, so we only report results for a one standard-error positive shock. The GIRFs are scaled so that the instantaneous response of the ACI index is $0.3$, making the responses comparable.
@@ 4.3 Impulse response analysis
Figure 2: Generalized impulse response functions to the identified ACI shock $h=0,1,....,36$ months ahead. From top to bottom, the responses of the ACI index, GDP growth rate, inflation growth rate, the interest rate variable, and transition weights of Regime 1 are depicted in the rows. The left column shows the responses to a positive one-standard-error shock with initial values generated from the stationary distribution of Regime 1. The right column shows the corresponding responses in Regime 2. All GIRFs have been scaled so that the instantaneous change in the ACI index is $0.3$. The $95\%$ confidence intervals are depicted with shaded grey areas. The results are based on $2500$ draws of initial values (for details of the algorithm, see Appendix B).
@@
Figure 2 presents the GIRFs (and the 95% confidence intervals) of the variables, including the transition weights, to the identified ACI shock $h=0,1,...,36$ quarters ahead. The right and left columns depict the responses to a one-standard-error positive shock in Regime 1 and Regime 2, respectively. In both regimes, a positive ACI shock mainly decreases GDP, consumer prices, and the interest rate. However, there is a major difference in the GIRFs between the regimes: the effects of the ACI shock on all macroeconomic variables are clearly stronger in Regime 1. This may be the case because this regime appears to be attributed to more volatile times, including recessions and crisis periods. Thus, it might represent a state of the economy that is more vulnerable to weather related shocks, and therefore, a similar shock is likely to result in greater damage in Regime 1 than in Regime 2. Other differences between the regimes are small.
The 95% confidence intervals depicted in Figure 2 convey the uncertainty about the initial state of the economy within each regime, as explained in Section 2.2. Hence, the very narrow confidence intervals in Regime 1 indicate only very weak dependence of the GIRFs on the initial values in that regime. This is because the ACI shock has almost no effect on the transition weights of Regime 1, as shown in the bottom left panel of Figure 2, and thereby the GIRFs resemble linear IRFs computed for this regime. In Regime 2, on the other hand, the confidence intervals are wider, implying some variation in the GIRFs depending on the initial values within the regime.
@@
As Regime 1 dominates the earlier part and Regime 2 is predominant in the latter part of the sample, we can conclude that the effect of severe weather shocks has diminished over time. This suggests that the U.S. economy has adapted to the effects of climate change. In contrast, Kim et al. (2022) found the effects of the ACI shock stronger at the end than at the beginning of their sample period from 1963:4 to 2019:5, which they interpret as insufficient adaptation in the U.S. economy to the changing distribution of weather related shocks. [^1] To check the robustness of our findings to the sample period, we also considered the subsample from 1963:4 to 2019:5 studied by Kim et al. (2022) and found the regimes of the fitted GSTVAR model quite similar to our main specification. Also, the estimated GIRFs and main conclusions from them are similar, with the exception that the interest rate variable increases in response to a positive ACI shock in Regime 1. The details and figures of the transition weights and GIRFs are presented in Appendix C.2. While our results are thus, in general, contradictory to theirs, there are some short periods in the latter part of our sample where Regime 1 with stronger impact of the ACI shock prevails. These particularly overlap with the major crises, which suggests that the potential adaptation may not provide sufficient resilience against weather related shocks in a turbulent state of the economy.
To evaluate the relative importance of the ACI shock in each regime, we compute the GFEVD for each regime separately based on the estimated generalized impulse responses. The relative contributions of the positive one-standard-error ACI shock to the forecast error variance at various horizons are presented in Table 2. The ACI shock dominates the ACI index, while the relative importance of the other shocks increases at longer horizons, especially in Regime 1. However, its importance for the macroeconomic variables is minor. In Regime 1, its relative contribution to movements in GDP growth is smaller than in Regime 2, which may follow from the fact that in this more turbulent regime, characterized by recessions and crises, presumably other shocks prevail that are more dominant drivers of the macroeconomy. The last two rows of Table 2 show that the relative importance of the ACI shock for the transition weights is small in Regime 1 but quite substantial in Regime 2. This suggests that the severe weather shocks may have some importance in the state of economy when the economy is in the more tranquil Regime 2, while the transition from the more turbulent Regime 1 towards Regime 2 is mainly driven by other shocks.
@@
ACI Regime 1 $100.00\%$ $91.54\%$ $87.24\%$ $80.82\%$ $69.76\%$ $62.27\%$
Regime 2 $100.00\%$ $98.96\%$ $96.82\%$ $95.80\%$ $93.45\%$ $91.10\%$
GDP Regime 1 $0.16\%$ $0.44\%$ $0.69\%$ $0.71\%$ $0.73\%$ $0.74\%$
Regime 2 $0.10\%$ $1.10\%$ $2.47\%$ $3.95\%$ $5.00\%$ $5.29\%$
CPI Regime 1 $5.40\%$ $3.58\%$ $3.24\%$ $3.07\%$ $3.01\%$ $3.00\%$
Regime 2 $0.04\%$ $0.59\%$ $0.65\%$ $0.75\%$ $0.96\%$ $1.06\%$
RATE Regime 1 $1.93\%$ $1.18\%$ $1.14\%$ $1.61\%$ $2.14\%$ $2.38\%$
Regime 2 $0.08\%$ $0.03\%$ $0.05\%$ $0.66\%$ $2.88\%$ $4.46\%$
$\boldsymbol{\alpha_{1,t}}$ Regime 1 $0.31\%$ $\phantom{1}0.80\%$ $\phantom{2}1.36\%$ $\phantom{2}1.48\%$ $\phantom{2}1.26\%$
Regime 2 $3.32\%$ $13.45\%$ $23.88\%$ $29.69\%$ $29.40\%$
@@
Table 2: The relative contribution of the ACI shock to the forecast error variances of the variables in each regime at the horizons $h=0,3,6,12,24,36$ based on the generalized forecast error variance decompositions in each regime. The relative contribution to the transition weights $\alpha_{1,t}$ is not defined (and thereby not shown) at impact, because the transition weights are $\mathcal{F}_{t-1}$-measurable.
@@ 5 Conclusion
We have introduced a Gaussian smooth transition vector autoregressive (GSTVAR) model with transition weights adopted from the Gaussian mixture VAR (GMVAR) model of Kalliovirta et al. (2016). Specifically, the transition weights of the GSTVAR model are constructed so that the greater the relative weighted likelihood of a regime is, the greater its transition weight is, which facilitates associating statistic characteristics and economic interpretations to the regimes. In other words, the GSTVAR model allows to capture gradual shifts in the dynamics of the data with switching dynamics that, for a $p$th order model, depend of the full distribution of the preceding $p$ observations.
@@
By making use of the results of Saikkonen (2008) and Kheifets and Saikkonen (2020), we have presented a sufficient condition for stationarity and ergodicity of the GSTVAR model. We have also discussed estimation of the model parameters by the method of maximum likelihood, which, due the complexity of the log-likelihood function, can be very tedious in practice. Hence, following Meitz et al. (2023) and Virolainen (2022, forthcoming), we propose employing a two-phase procedure in which a genetic algorithm is used to find starting values for a gradient based method. The introduced methods, including a modified version of a genetic algorithm, are implemented in the CRAN distributed R package sstvars (Virolainen, 2024). A potential area of future research is an extension of our model that accommodates strong forms of conditional heteroskedasticity by incorporating autoregressive conditional heteroskesticity or stochastic volatility to the shocks.
@@
We have applied our method to study the macroeconomic effects of severe weather shocks. Our monthly data set covers the period from 1961:1 to 2022:3 and contains an indicator of the frequency of severe weather and the extent of sea level rise in addition to a number of macroeconomic variables. Our structural GSTVAR model has two regimes. Regime 1 mainly dominates the earlier part of the sample period, particularly the turbulent times of 1970’s and 1980’s, but also prevails in the later sample during the Financial crisis and the COVID-19 crisis, while the Regime 2 dominates the latter part of the sample period. Following Kim et al. (2022), we identified the severe weather shock recursively, placing the severe weather indicator first in the vector of variables. A positive weather shock was found to decrease GDP, consumer prices, and the interest rate in both regimes, but the effects are stronger in Regime 1. This finding contradicts the main result of Kim et al. (2022), who found the impacts of the weather shock to get stronger over time, interpreting this as lack of adaptation to changing weather. In contrast, our findings suggest that the U.S. economy has adapted to the changing distribution of weather related shocks. However, the adaptation may not provide sufficient resilience when the economy is in a vulnerable state.
@@ Appendix B Monte Carlo algorithm for estimating the GIRF
We present a Monte Carlo algorithm for estimating the generalized impulse response function defined in Equation (2.6) for initial values $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ Our algorithm is adapted from Koop et al. (1996, pp. 135-136) and Kilian and Lütkepohl (2017, pp. 601-602). We assume that the history $\boldsymbol{y}_{t-1}$ follows a known distribution $G$, which may be such that it produces a single outcome with probability one (corresponding to a fixed $\boldsymbol{y}_{t-1}$), or it can be drawn from a specific regime as described in Section (2.2), for instance. In the following, $y_{t+h}^{(i)}(\delta_{j},\boldsymbol{y}_{t-1})$ denotes a realization of the process at time $t+h$ conditional on the structural shock of sign and size $\delta_{j}$ in the $j$th element of $e_{t}$ hitting the system at time $t$ and on the $p$ observations $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ preceding the time $t$, whereas $y_{t+h}^{(i)}(\boldsymbol{y}_{t-1})$ denotes an alternative realization conditional on the history $\boldsymbol{y}_{t-1}$ only.
The algorithm proceeds with the following steps.
@@
    Set the horizon $H$, the numbers of repetitions $R_{1}$ and $R_{2}$, and the sign and size $\delta_{j}$ for the $j$th structural shock that is of interest.
@@
    Use the modified reduced form shock $u_{t}^{*}$ and the rest $H$ structural shocks $\varepsilon_{t}$ obtained from Step 2 to compute realizations $y_{t+h}^{(i)}(\delta_{j},\boldsymbol{y}_{t-1})$ for $n=0,1,...,H$, iterating forward. At $h=0$, the modified reduced form shock $u_{t}^{*}$ calculated from the structural shock $e_{t}^{*}$ in Step 2 is used. From $h=1$ onwards, the $h+1$th structural shock $e_{t}$ is used to calculate the reduced form shock $u_{t+h}=B_{t+h}e_{t+h}$.
@@
    Use the rest $H+1$ the structural shocks $e_{t}$ obtained from Step 2 to compute realizations $y_{t+h}^{(i)}(\boldsymbol{y}_{t-1})$ for $h=0,1,...,H$, so that the non-modified reduced form shock $u_{t}=B_{t}e_{t}$ is used to compute the time $h=0$ realization. Otherwise proceed similarly to the previous step.

### v3

@@ 4.3 Structural analysis
We are predominantly interested in the economic impact of severe weather, and, therefore, a severe weather shock (ACI shock) must be identified. This amounts to imposing a unique structure of the impact matrix $B_{t}$ in (2.3) governing the contemporaneous relationships of the shocks, so that one of the shocks can be labelled the ACI shock. Following Kim et al. (in press), our key identification restriction is that the ACI shock is the only shock that can instantaneously affect the ACI. This identification restriction seems reasonable, as it states that the other (macroeconomic) shocks do not affect severe weather or sea level within a month, but allows them to affect the ACI in the long run. We establish the identification by placing the ACI first in the vector of variables (with the rest of the variables in the order GDP, CPI, RATE) and imposing a recursive lower-triangular structure on the impact matrix $B_{t}$, obtained by the Cholesky decomposition of the estimated conditional error covariance matrix $\Omega_{y,t}$.
To study the effects of the ACI shock, we compute the generalized impulse response functions of the variables to it. As pointed out in Section 2.2, since the transition weights are endogenously determined and the regime can shift as a result of a shock, the impulse responses generally depend on the initial values as well as on the sign and size of the shock. We are particularly interested in finding out about the potential state-dependence of the effects of the ACI shock. To capture them, we take all such histories of length $p$ from the data that indicate the dominance of a given regime, and compute the GIRFs for each regime separately. Specifically, for Regime $\tilde{m}$, we take the histories $\boldsymbol{y}_{t-1}$ for which the corresponding transition weight $\alpha_{\tilde{m},t}$ is greater than $0.75$. With the threshold $0.75$, the given regime is clearly dominant but a reasonably large amount of histories are still included in both regimes (Regimes 1 and 2 have $115$ and $603$ such histories, respectively). To closely match the properties of the data, for each history, the GIRF to the corresponding recovered structural shock is computed. Finally, the GIRFs are scaled so that the instantaneous response of the ACI is $0.3$, making the responses comparable.
@@
Figure 2: Generalized impulse response functions to the identified ACI shock $h=0,1,....,36$ months ahead. From top to bottom, the responses of ACI, GDP growth rate, CPI growth rate, the interest rate variable, and transition weights of Regime 1 are depicted in the rows. The left and right columns show the responses conditional on Regime 1 and Regime 2 dominating when the shock arrives, respectively. All GIRFs have been scaled so that the instantaneous change in the ACI index is $0.3$. The results are based on $2500$ Monte Carlo repetitions for each history (for details of the algorithm, see Appendix B).
Figure 2 presents the sample distribution (over the histories) of the GIRFs of the variables, including the transition weights, to the identified ACI shock $h=0,1,...,36$ quarters ahead. Each light gray curve with low opacity depicts the GIRF corresponding to one history $\boldsymbol{y}_{t-1}$, so the darkness of a region in the figure indicates a greater concentration of GIRFs in that area. The left and right columns depict the responses in Regime 1 and Regime 2, respectively. In Regime 1, the vast majority of the GIRFs follow virtually the same sample paths. This is the case because the ACI shock has almost no effect on the transition weights of Regime 1, as shown in the bottom left panel. Although the majority of the GIRFs follow similar sample paths also in Regime 2, they exhibit more variation depending on the history $\boldsymbol{y}_{t-1}$ (and the related sign and size of the shock).
@@
In both regimes, a positive ACI shock mainly decreases GDP, consumer prices, and the interest rate. However, there is a major difference in the GIRFs between the regimes: the effects of the ACI shock on all macroeconomic variables are clearly stronger in Regime 1. This may be the case because Regime 1 appears to be attributed to more volatile times, including recessions and crisis periods. Thus, it might represent a state of the economy that is more vulnerable to weather related shocks, and therefore, a similar shock is likely to result in greater damage in Regime 1 than in Regime 2. Other differences between the regimes are minor.
@@
As Regime 1 dominates the earlier part and Regime 2 is predominant in the latter part of the sample period, it seems reasonable to conclude that the effect of severe weather shocks has diminished over time. This suggests that the U.S. economy has adapted to the effects of climate change. In contrast, Kim et al. (in press) found the effects of the ACI shock stronger at the end than at the beginning of their sample period from 1963:4 to 2019:5, which they interpreted as insufficient adaptation of the U.S. economy to the changing distribution of weather related shocks. While our results are thus, in general, contradictory to theirs, there are some short periods in the latter part of our sample where Regime 1 with stronger impact of the ACI shock prevails. These particularly overlap with the major crises, which suggests that the potential adaptation may not provide sufficient resilience against weather related shocks in a turbulent state of the economy.
@@
ACI All data $100.00\%$ $85.73\%$ $80.32\%$ $77.22\%$ $72.69\%$ $69.73\%$
Regime 1 $100.00\%$ $73.04\%$ $67.88\%$ $62.06\%$ $55.24\%$ $51.50\%$
Regime 2 $100.00\%$ $88.25\%$ $82.81\%$ $80.39\%$ $76.44\%$ $73.44\%$
GDP All data $5.38\%$ $10.60\%$ $14.34\%$ $16.74\%$ $17.57\%$ $17.39\%$
Regime 1 $5.96\%$ $\phantom{1}4.49\%$ $\phantom{1}6.37\%$ $\phantom{1}6.12\%$ $\phantom{1}5.93\%$ $\phantom{1}5.77\%$
Regime 2 $5.25\%$ $11.83\%$ $16.01\%$ $18.97\%$ $20.03\%$ $19.85\%$
CPI All data $\phantom{1}5.84\%$ $\phantom{1}7.70\%$ $7.45\%$ $7.25\%$ $7.43\%$ $7.47\%$
Regime 1 $18.92\%$ $11.01\%$ $9.93\%$ $8.94\%$ $8.19\%$ $7.86\%$
Regime 2 $\phantom{1}3.39\%$ $\phantom{1}7.17\%$ $7.07\%$ $7.02\%$ $7.36\%$ $7.50\%$
RATE All data $\phantom{1}3.53\%$ $2.21\%$ $2.21\%$ $5.72\%$ $11.88\%$ $14.58\%$
Regime 1 $15.30\%$ $9.66\%$ $8.66\%$ $10.23\%$ $11.90\%$ $12.30\%$
Regime 2 $\phantom{1}1.14\%$ $0.73\%$ $0.97\%$ $4.92\%$ $12.05\%$ $15.21\%$
$\boldsymbol{\alpha_{1,t}}$ All data $13.53\%$ $13.77\%$ $14.66\%$ $14.71\%$ $13.72\%$
Regime 1 $\phantom{1}5.55\%$ $\phantom{1}5.20\%$ $\phantom{1}4.51\%$ $\phantom{1}4.49\%$ $\phantom{1}4.71\%$
Regime 2 $15.24\%$ $15.60\%$ $16.81\%$ $16.85\%$ $15.62\%$
@@
Table 2: The relative contribution of the ACI shock to the forecast error variances of the variables in each regime at the horizons $h=0,3,6,12,24,36$ based on the generalized forecast error variance decompositions calculated using the length $p$ histories of the data with the related sizes and signs of the shock. The rows labeled ”All data” present the GFEVDs based on GIRFs using all the length $p$ histories of the data with the relates signs and sizes of the shock, whereas rows labeled ”Regime 1” or ”Regime 2” shows the GFEVDs based on only such length $p$ histories that indicate the dominance of the specified regime (i.e., the its related transition weight is larger than $0.75$). The relative contribution to the transition weights $\alpha_{1,t}$ is not defined (and thereby not shown) at impact, because the transition weights are $\mathcal{F}_{t-1}$-measurable.
To assess the relative importance (or economic significance) of the ACI shock, we compute the generalized forecast error variance decompositions using all the histories $\boldsymbol{y}_{t-1}$, $t=1,...,T$, in the data as the initial values with the corresponding sign and size of the recovered structural shock. The rows labeled “All data” in Table 2 present the relative contribution of the ACI shock to the forecast error variance of the variable in question at each horizon, while the rows labeled “Regime 1” and “Regime 2” contain the corresponding figures for each regime (i.e., for the histories with a transition weight of the regime in question higher than $0.75$ in time period $t$). The ACI shock dominates the ACI index, but the other shocks become relatively more important at longer horizons. For the remaining variables, the ACI shock plays a lesser, yet clearly notable, role, contributing approximately 5–17%, 6–8% and 2–15% to the forecast error variance of GDP, CPI and the interest rate in the entire sample period, respectively. The ACI shock also seems to have some importance for regime shifts, with its relative contribution to the forecast error variance of the transition weight $\alpha_{1,t}$ hovering between 13 and 15%. The ACI shock is economically significant also in both regimes. Compared to Regime 1, its relative importance is, in general, greater for GDP and the transition weight in Regime 2, whereas the opposite holds for CPI and the interest rate, especially at short horizons.
4.4 Robustness checks
To check the robustness of our findings with respect to the sample period, we also consider the subsample period from 1963:4 to 2019:5 considered by Kim et al. (in press). The results general conclusion remain intact in that the regimes of the fitted GSTVAR model quite similar to our main specification. Also, the estimated GIRFs and are similar, with the exception that the interest rate variable increases in response to a positive ACI shock in Regime 1. The details, including the evolution of the transition weights and the GIRFs can be found in Appendix E.2.1.
For comparison, we also report the results of the GMVAR model (Kalliovirta et al., 2016) as well as the recursively identified linear Gaussian SVAR model (obtained as special case of GMVAR and GSTVAR models with $M=1$) for the entire sample period. The mixing weights of the fitted GMVAR model, presented in Figure E.7 in Appendix E.2.2 are rather similar to the transition weights of our GSTVAR model, as expected due to the similar function forms of the weight functions. Also the GIRFs, presented in Figure E.8 in Appendix E.2.2, are quite similar to those in our GSTVAR model. In particular, in line with our GSTVAR model, the GMVAR model displays clearly stronger responses of GDP, inflation and the interest rate to the ACI shock in the more turbulent Regime 1 than in Regime 2. Finally, the finding that a positive ACI shock causes a decrease in all GDP, inflation and the interest is quite robust, as it emerges in all models considered, including the recursive Gaussian SVAR model, whose impulse responses are also depicted in Figure E.8 in Appendix E.2.2.
@@ 5 Conclusion
We have introduced a Gaussian smooth transition vector autoregressive (GSTVAR) model with transition weights adopted from the Gaussian mixture VAR (GMVAR) model of Kalliovirta et al. (2016). However, compared to the GMVAR model, our model has the advantage of great flexibility in that it enables capturing gradual shifts in the dynamics of the data, whereas the GMVAR model involves discrete regimes. On the other hand, switching between the regimes in the $pt$h-order GSTVAR model depends on the full distribution of the preceding $p$ observations, in contrast to previously introduced other smooth transition VAR models, in which it is governed by specific switching variables. Hence, our model facilitates associating the regimes with statistical information in the data in a versatile manner. The structural counterpart of the GSTVAR model has the desirable feature that it produces the dynamic effects of the structural shocks on the transition weights governing regime switching.
@@
We have discussed estimation, model selection and diagnostic checking as well as conducting structural analysis in the GSTVAR model. Moreover, by making use of the results of Saikkonen (2008) and Kheifets and Saikkonen (2020), we have introduced a sufficient condition for stationarity and ergodicity of the model. Due to the complexity of the log-likelihood function, estimation by the method of maximum likelihood can be quite tedious in practice, and following Virolainen (2022, 2025b) (and others), we propose a two-phase procedure in which a genetic algorithm is used to find starting values for a gradient based estimation method. The introduced methods, including a modified version of a genetic algorithm, are implemented in the CRAN distributed R package sstvars (Virolainen, 2025a). A potential area of future research is an extension of our model that accommodates conditional heteroskedasticity by incorporating autoregressive conditional heteroskesticity or stochastic volatility to the shocks.
@@
We have applied our method to study the macroeconomic effects of severe weather shocks in the U.S. Our monthly data set covers the period from 1961:1 to 2022:3 and contains an indicator of the frequency of severe weather and the extent of sea level rise in addition to a number of macroeconomic variables. Our structural GSTVAR model has two regimes, one of which (Regime 2) prevails in the latter part of the sample period, while the other regime (Regime 1) mainly dominates its earlier part, particularly the turbulent times of 1970s and 1980s, but also prevails in the later sample during the Financial crisis and the COVID-19 crisis. Kim et al. (in press) have recently also studied the economic impact of severe weather in the U.S., and following their lead, we have identified the severe weather shock recursively, placing the severe weather indicator first in the vector of variables. A positive weather shock was found to decrease GDP, consumer prices, and the interest rate in both regimes, but the effects are stronger in Regime 1. Hence, our results suggest that the U.S. economy, with the exception of certain crisis periods, has adapted to the changing distribution of weather related shocks. This finding contradicts the main result of Kim et al. (in press), who found the impact of the weather shock to get stronger over time, interpreting this as lack of adaptation to changing weather.
@@ Appendix B Monte Carlo algorithm for estimating the GIRF
We present a Monte Carlo algorithm for estimating the generalized impulse response function defined in Equation (2.6) for initial values $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$. Our algorithm is adapted from Koop et al. (1996, pp. 135-136) and Kilian and Lütkepohl (2017, pp. 601-602). We assume that one is either interested on a known set $S$ of the histories $\boldsymbol{y}_{t-1}$ (e.g., certain histories or a history taken from the data), or the history $\boldsymbol{y}_{t-1}$ follows a known distribution $G$ (e.g., the stationary distribution of a specific regime). In the following, $y_{t+h}^{(i)}(\delta_{j},\boldsymbol{y}_{t-1})$ denotes a realization of the process at time $t+h$ conditional on the structural shock of sign and size $\delta_{j}$ in the $j$th element of $e_{t}$ hitting the system at time $t$ and on the $p$ observations $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ preceding the time $t$, whereas $y_{t+h}^{(i)}(\boldsymbol{y}_{t-1})$ denotes an alternative realization conditional on the history $\boldsymbol{y}_{t-1}$ only.
For a single history $\boldsymbol{y}_{t-1}$ and given sign and size $\delta_{j}$ of the $j$th structural shock (that is of interest), the GIRF can be estimated with the following steps.
@@
    Set the horizon $H$ and the number of repetitions $R_{1}$.
@@
    Use the modified reduced form shock $u_{t}^{*}$ and the rest $H$ structural shocks $\varepsilon_{t}$ obtained from Step 1 to compute realizations $y_{t+h}^{(i)}(\delta_{j},\boldsymbol{y}_{t-1})$ for $n=0,1,...,H$, iterating forward. At $h=0$, the modified reduced form shock $u_{t}^{*}$ calculated from the structural shock $e_{t}^{*}$ in Step 1 is used. From $h=1$ onwards, the $h+1$th structural shock $e_{t}$ is used to calculate the reduced form shock $u_{t+h}=B_{t+h}e_{t+h}$.
@@
    Use the rest $H+1$ the structural shocks $e_{t}$ obtained from Step 1 to compute realizations $y_{t+h}^{(i)}(\boldsymbol{y}_{t-1})$ for $h=0,1,...,H$, so that the non-modified reduced form shock $u_{t}=B_{t}e_{t}$ is used to compute the time $h=0$ realization. Otherwise proceed similarly to the previous step.